OpenStudy (anonymous):

Solve 3x + 2y = -12 for y. give a medal to first that answers

OpenStudy (amoodarya):

3x+2y=-12 2y=-12-3x divide by 2 y=-6=3/2 x

OpenStudy (anonymous):

ah thank you

OpenStudy (anonymous):

a medal for you sir